OpenStudy (anonymous):

3. 7x^2 – 144 = –x^4 (1 point) please help, & walk thru...

OpenStudy (anonymous):

3, –3, 4i, 0 3, 4i 3, –3, 4i, –4i –3, –4i

OpenStudy (anonymous):

those are the answer choices, but i'd like to learn how to solve this too.

OpenStudy (anonymous):

\[7x^2 - 144 = -x^4\] \[x^4 + 7x^2 - 144 = 0\] let p = x^2, then: \[p^2 + 7p - 144 = 0\] This is just a quadratic equation which you can solve for p, and then use p = x^2 to find out what x can be

OpenStudy (anonymous):

so I added x^4 to each side?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

And Im confused on what to do after that?

OpenStudy (anonymous):

so then do i add 144 to each side?

OpenStudy (anonymous):

No, you let p = x^2 and substitute that into the original equation. Then you get a quadratic equation you can solve by using the quadratic formula

OpenStudy (anonymous):

Ohhh so it's C? 3,-3, 4i, -4i?

OpenStudy (anonymous):

Thank u :-) Got it right!